– General Linux 1 – Configure communication devices [2]

(Linux Professional Institute Certification)

a

   .~.
   /V\      by: geoffrey robertson
  // \\         geoffrey@zip.com.au
  @._.@
$Id: gl1.101.6.slides.tex,v 1.1 2003/02/19 02:03:06 geoffr Exp $

aCopyright \begin{matrix}c \\ ◯ \end{matrix} 2002 Geoffrey Robertson. Permission is granted to make and distribute verbatim copies or modified versions of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.

Configure communication devices [2]

(1.1) 101 Hardware & Architecture [11]

1.101.1 Configure BIOS Fundamental Settings [3] 1.101.3 Configure Modem and Sound cards [2] 1.101.4 Setup SCSI devices [2] 1.101.5 Setup different PC expansion cards [2] 1.101.6

Configure Communication Devices [2] 1.101.7 Set Up USB Devices [2]

Configure communication devices [2]

Objective

Configure communication devices [2]

Key files, terms, and utilities

Configure communication devices [2]

Resources of interest The End

6